# Model Provenance Testing for Large Language Models

### v1

@@ Abstract
Large language models are increasingly customized through fine-tuning and other adaptations, creating challenges in enforcing licensing terms and managing downstream impacts. Tracking model origins is crucial both for protecting intellectual property and for identifying derived models when biases or vulnerabilities are discovered in foundation models. We address this challenge by developing a framework for testing model provenance: Whether one model is derived from another. Our approach is based on the key observation that real-world model derivations preserve significant similarities in model outputs that can be detected through statistical analysis. Using only black-box access to models, we employ multiple hypothesis testing to compare model similarities against a baseline established by unrelated models. On two comprehensive real-world benchmarks spanning models from 30M to 4B parameters and comprising over $600$ models, our tester achieves $90-95$% precision and $80-90$% recall in identifying derived models. These results demonstrate the viability of systematic provenance verification in production environments even when only API access is available.
@@ 1 Introduction
Platforms such as Amazon SageMaker and Hugging Face have enabled wide scale distribution of ML models, most notably large language models (LLMs) [40, 50, 22]. Certain models, called foundational models, require extensive computational resources and datasets to train. For instance, it was reported in [1] that Meta used two $24,000$ GPU clusters to train Llama 3. But foundational models are relatively cheap to fine-tune or customize for downstream applications [25, 51]. We are thus seeing a proliferation of fine-tuned models.
The increase of publicly available foundation models and datasets, however, has also triggered concerns over unauthorized use of intellectual property (IP) and concerns of compromised datasets [10] or models [20]. These issues are present not just in open-source ecosystems, but also for proprietary models that are hidden behind APIs [4, 33]. For instance, concerns about model stealing attacks wherein one can extract the model parameters even for production-level models are on the rise [11, 47, 34]. Similarly, there is a growing concern that proprietary models may contain backdoors or vulnerabilities, making them susceptible to jailbreaking [20, 3, 56]. Despite best efforts to create a safe environment for the development of (public and commercial) foundation models, there have already been instances of reported misuse [38, 18, 46].
This landscape highlights the growing need for model provenance testing. The problem of model provenance testing is as follows: Determine whether a target model has been derived from a foundational model by lighter customization such as fine-tuning. This problem has applications in tracking reuse of models not just in open marketplaces but also across product teams in large organizations. Current policy regulations such as GDPR [14] and Artificial Intelligence (AI) Safety Act [13] require security and privacy compliance for all AI-enabled apps [16]. When a security or privacy audit finds a problem with a foundational model, it becomes important to identify which other models in use by the organization may be derived from the problematic one and take appropriate remedial actions (e.g. revoke, retrain, or fortify) to mitigate the risk of further non-compliant use. Model provenance tracking is often useful after the fine-tuned model has been deployed, and when authentic ground truth is unavailable or unreliable.
@@
In this paper, we design the first practical model provenance tester for LLMs that requires only query access. Our proposed techniques stem from a key empirical observation: The output distribution of fine-tuned models is often close to that of their parent model. This distance between a model and its true parent is typically smaller than that between the model and another unrelated models, making it possible to reliably trace back a derived model to the original parent. In order to keep assumptions to a minimum, we propose to employ the principled framework of statistical hypothesis testing. Specifically, we use black-box sampling and estimation to determine whether the distribution captured by the given model is close to that of the candidate parent. Such estimation can provide formal statistical significance measures, which can be used to check for the null hypothesis, i.e., the customized LLM is not close to the given parent model.
We conduct an extensive empirical evaluation across two comprehensive benchmarks comprising over $600$ models from Hugging Face, ranging from $30$M to $4$B model parameters. Our tester achieves $90-95$% precision and $80-90$% recall in detecting model provenance relationships, even with a limited number of queries. Therefore, while being simple, we find that our proposed method achieves excellent accuracy.
Contributions. We present the problem of model provenance testing. Our work initiates the study of such testers in the context of customized LLMs, and keeps assumptions minimal, i.e., having only black-box query access to models. We show that in the existing landscape of open-source LLM models, it is practical to fairly accurately determine provenance.
2 Overview
Model provenance testing has many motivating applications, but we present one representative scenario for concreteness.
2.1 Motivating Scenario
Pretraining large language models (LLMs) involves significant investment, requiring substantial computational resources costing millions of dollars in infrastructure and thousands of GPU hours. When Company A releases a pretrained LLM (called LLM-$f$), it employs specific licensing terms crucial for protecting this investment, maintaining competitive advantage, and controlling the model’s usage [33, 4]. These licenses typically include restrictions on commercial usage, model modification, or redistribution, and may incorporate provisions for monitoring downstream usage and revenue sharing.
Startup B might download LLM-$f$, perform only fine-tuning, but claim to have pretrained their model (LLM-$g$) from scratch, thereby circumventing licensing requirements and misrepresenting company A’s work. In such cases, we want to be able to determine if LLM-$g$ is derived through fine-tuning LLM-$f$ and resolve the model provenance problem. Given the significant economic incentive to fine-tune existing models rather than pretrain from scratch, verifying model provenance becomes essential for enforcing licensing terms and protecting intellectual property rights. Aside from fine-tuning, there are a few other inexpensive techniques (mixture-of-experts, prompt engineering) that Startup B could employ to derive LLM-$g$ from LLM-$f$ while still claiming independent development. All these approaches require orders of magnitude less computational resources than pretraining, making them particularly attractive for circumventing licensing restrictions. Thus, the model provenance problem extends beyond detecting fine-tuning to identifying any form of model derivation that might violate licensing terms. We have witnessed such a case recently when the pretrained model Llama 3.1 was reportedly used for military operations, violating the license agreement [38].
Startup B can release LLM-$g$ either by making its weights publicly available (e.g., on Hugging Face) or by offering it as an online service through an API (e.g., ChatGPT or PaLM-2). The API-like interface is more general than if weights were also available to a tester trying to determine provenance. In this setup, the tester must rely solely on the model’s responses to input queries. This distinction significantly impacts the techniques available for detecting whether LLM-$g$ is derived from LLM-$f$.
@@ 2.1 Motivating Scenario
We consider a testing framework with minimal assumptions only, where the tester can query LLM-$g$ on arbitrary prompts (through the API) and get responses. The tester has no access to the training datasets, hyperparameters used by either company, or any information about potential modifications performed by Startup B. This mirrors real-world conditions where companies do not always disclose their training procedures, data sources, or modification techniques, making the provenance testing problem both practical and challenging.
2.2 Our Approach
@@
Our approach to testing model provenance is based on a key observation: fine-tuning and other model derivation techniques typically result in only limited changes to the original model, as they primarily adapt the model for new tasks. After fine-tuning, the derived model LLM-$g$ may remain similar to its parent model LLM-$f$, as the process focuses on refining specific capabilities rather than creating fundamental changes to the model distribution.
In our setting, the tester can only interact with models through their API interfaces, submitting prompts and analyzing the generated outputs. We thus compute similarity of models by analyzing the behavioral patterns between LLM-$f$ and LLM-$g$ through their responses. For efficiency and simplicity, we focus on comparing next-token predictions on random prompts rather than analyzing entire sequences. While extending the analysis to longer sequences ($n$-grams) might provide additional signals, our empirical evidence suggests that the next-token distribution already captures significant information about a model’s decision-making patterns and internal representations.
Our method tests a number of random diverse prompts, looking for evidence that LLM-$g$’s output distributions consistently align more closely with LLM-$f$ than would be expected from a model that was not derived from LLM-$f$. However, there is no fixed threshold for what constitutes “close alignment”. We cannot simply say that matching on 10% or 20% of outputs indicates provenance, as the degree of similarity can vary significantly depending on the intensity of fine-tuning, how different the downstream task is and other modifications performed. Even models from which LLM-$g$ was not derived may occasionally produce matching outputs, and we need to understand this rate of coincidental alignment. We thus need more rigorous and universal approach, that is based not only on similarity between models, but on finding a level of similarity that is unusually higher than coincidence.
To establish a meaningful baseline, i.e., to characterize the expected level of coincidental agreement between unrelated models, we take several additional control models and analyze how closely their outputs align with LLM-$g$. The selection of control models is crucial for establishing reliable baselines for coincidental agreement. These control models are selected to represent a diverse range of output distributions, thus allowing to better estimate the average levels of similarity that might occur between unrelated models. If the similarity between LLM-$f$ and LLM-$g$ is substantially higher than the similarity observed with the control models, it provides compelling evidence that LLM-$g$ was indeed derived from LLM-$f$, as their behavioral patterns exhibit a level of consistency that cannot be easily attributed to random chance. Conversely, if the similarity falls within the typical range observed among the control models, it may indicate that the observed alignment is not sufficient to conclusively determine provenance.
The actual test for excess similarity between LLM-$f$ and LLM-$g$ is performed using multiple hypothesis testing. Each hypothesis is formulated to show that similarity of LLM-$g$ to LLM-$f$ is greater than the similarity to any control models. When all these hypothesis tests yield statistically significant results, we obtain strong evidence that the similarity between these models is systematically higher than what would be expected by chance. This statistical framework offers several key advantages. First, it handles varying degrees of fine-tuning adaptively, without requiring assumptions about the expected similarity between LLM-$g$ and LLM-$f$. The test naturally identifies any significant bias in matching rates compared to the best non-parent baseline model. Second, rather than relying on arbitrary thresholds, it provides a principled way to quantify confidence in provenance determination through $p$-values, i.e., observed bias in matching rates, and confidence intervals. Third, the approach is computationally efficient, requiring only a modest number of test prompts to achieve statistical significance.
2.3 Advantages and Effectiveness
Our approach stands out in its reliance on minimal assumptions, particularly the requirement of only black-box query access to the models in question. By operating without access to proprietary model weights, training data, or hyperparameters, we mirror real-world conditions where such internal details are typically inaccessible. This makes our method highly practical and broadly applicable, as it can be employed in scenarios where models are available exclusively through API endpoints with restricted internal information.
A key strength of our method is its statistical foundation. By employing multiple hypothesis testing, we objectively assess whether the observed similarities between LLM-$g$ and LLM-$f$ are significant beyond what could be attributed to random chance. This framework allows us to compute precise $p$-values and confidence intervals, providing quantifiable evidence to support our conclusions about model similarities. Importantly, our approach avoids reliance on arbitrary thresholds and naturally adapts to varying degrees of similarity.
From a practical standpoint, our method is both efficient and scalable. Focusing on next-token predictions over a diverse set of random prompts minimizes the computational overhead and the number of queries required. This efficiency is particularly valuable when dealing with large models and the inherent limitations of API access, such as query costs and rate limits. The simplicity of our approach allows for straightforward implementation without the need for specialized hardware or extensive computational resources.
Regarding effectiveness, our method has demonstrated high accuracy in detecting model provenance. As we will detail in Section 4, our approach consistently identifies provenance pairs with over $90$% precision and $80$% recall when tested on large set of pairs. These results highlight the method’s capability to reliably determine provenance, even when the derived model has been significantly fine-tuned or altered. This effectiveness underscores its practical utility in real-world applications where accurate provenance detection is critical.
2.4 Challenges
While our approach offers a robust and practical solution to the model provenance problem, it is important to acknowledge its limitations and the challenges inherent in this domain. Understanding these factors is crucial for interpreting the results of our method and guiding future enhancements.
First, a primary challenge is dealing with cases where the derived model LLM-$g$ has undergone extensive modifications that significantly alter its behavior relative to the parent model LLM-$f$. Such modifications might include heavy fine-tuning on large and diverse datasets, architectural changes, or techniques designed to intentionally mask the model’s origins. In these scenarios, the residual similarities between LLM-$g$ and LLM-$f$ may diminish to the point where our statistical tests lose sensitivity, potentially leading to false negatives.
Second, our method depends on the selection of control models used for baseline comparisons. The effectiveness of our statistical tests hinges on having appropriate control models that are sufficiently diverse and representative of unrelated models. If the control models are inadvertently dissimilar to LLM-$f$, this could affect the baseline similarity levels and impact the test’s reliability. Careful consideration is therefore required in choosing control models to ensure they provide a meaningful contrast in the hypothesis testing framework.
Another challenge lies in the inherent variability of language models and the possibility of coincidental similarities. Language models trained on similar datasets might exhibit overlapping behaviors, even if one is not derived from the other. This could potentially lead to false positives, where our method incorrectly identifies a provenance relationship due to shared patterns that are common in the domain rather than indicative of direct derivation. Addressing this requires careful interpretation of the results and may need an additional analysis to rule out such coincidental similarities.
In Section 4 we will experimentally determine the impact of some of these assumptions on the tester’s effectiveness.
3 Main Approach
In this section we provide a more formal introduction of the problem and present the main model provenance tester.
3.1 Problem Setup
A large language model (LLM) takes as input a sequence of tokens $x_{<i}=(x_{1},\ldots,x_{i-1})$ from a fixed set of tokens $\mathcal{V}$ and outputs the next token $x_{i}\in\mathcal{V}$, i.e., $f_{\theta}(x_{<i})=x_{i}$. Formally, LLM is denoted with $f_{\theta}$, defined by its architecture $f$ and its model parameters $\theta$. Herein, it is sufficient for us to consider the LLM as a mapping $f_{\theta}:\mathcal{V}^{I}\rightarrow\mathcal{V}$, where $I$ is an arbitrary sequence length. Conditioned on an input sequence $x_{<i}$, the $f_{\theta}$ induces a probability distribution on the whole vocabulary set $\mathcal{V}$ so each token is assigned a probability score. In the sequel, we assume $f_{\theta}$ chooses the token with the highest probability, so-called temperature $t=0$ case. Note, when $f_{\theta}$ uses higher temperatures $t>0$, the most likely token can still be empirically determined by repeatedly sampling from the induced distribution and selecting the most frequently occurring token.
A model $f_{\theta}$ can be customized through various techniques to create a new model. The most common approach is fine-tuning, where the original parameters $\theta$ are updated using gradient descent on a new dataset to obtain $\theta^{\prime}$. Other customization techniques such as mixture-of-experts can involve creating a model $g_{\theta^{\prime}}$ with different architecture that reuses parts of the architecture of the initial model $f_{\theta}$. Thus, in the sequel, when addressing customizations and models, we omit parameter notations, and denote the initial (parent) model as $f$, and the derived (child) model as $g$, and say that $(f,g)$ constitutes a model provenance pair. Importantly, these customizations typically require orders of magnitude less computational resources than training from scratch. Under our setup, we have only query access to the models - we can provide input sequences $x_{<i}=(x_{1},\ldots,x_{i-1})$ and observe their corresponding output tokens $x_{i}$. This reflects real-world scenarios where models are accessed through APIs, while keeping internal model details private. The model provenance problem consists thus in detecting if $g$ was derived from $f$, given only query access to models.
Threat Model.
Our proposed methods are primarily evaluated for a non-adaptive adversary. The model customizer is not aware of the strategy of the model provenance tester. They are not aware of the knowledge the tester has, i.e., the control models, the set of randomly sampled inputs, or the parameters of the statistical test. On the other hand, one may consider an adaptive adversary, i.e., customizers that may employ techniques specifically designed to evade provenance detection, especially based on the knowledge of the provenance tester. For instance, they might introduce noise or deliberate alterations to the output distributions to reduce detectable similarities with the parent model. Such adaptive obfuscation pose a significant challenge, as they can diminish the effectiveness of our statistical tests. Developing methods to counteract such adaptive adversarial strategies is beyond the scope of this work.
3.2 Model Provenance Tester
We formulate the model provenance question as the problem of checking whether two models are similar. This approach is based on the observation that models derived from one another tend to exhibit higher similarity compared to unrelated models. We now detail how the similarity check is implemented.
Since the tester has only query access to the models, the only data it can collect is from providing inputs (called prompts) and analyzing the corresponding output tokens. Furthermore, due to our minimal assumptions and lack of information about the training datasets, the tester queries the models on randomly chosen short sequences designed to avoid predictable (low-entropy) continuations. For example, the tester avoids prompts like “To be or not to” since the next token is highly likely to be “be” a common continuation that many models would predict. [^1] Tokens with low entropy are those for which the next token can be easily predicted by any model, making them unsuitable for distinguishing between models. We further denote such prompt space with $\Omega$ and sample from it uniformly at random.
The tester independently samples from $\Omega$ a set of $T$ such prompts, queries each model on the same set of prompts, and then compares their output tokens pairwise. For each prompt, we compare only the first output token generated by each model; however, in principle, $n$-grams (sequences of $n$ consecutive tokens) could also be considered. The similarity between two models is then calculated as the proportion $\mu$ of prompts on which the models produce the same output token:
$\mu=\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(f(x_{<i}^{j})=g(x_{<i}^{j}))$
However, the ratio $\mu$ by itself may not be sufficient to determine whether $f$ and $g$ are truly similar. We need to understand if the similarity $\mu$ between $g$ and $f$ is higher than what we would expect if $g$ were unrelated to $f$. To make this assessment, we need to establish what level of similarity we should expect between $g$ and models that are unrelated to it but share key characteristics with $f$. We do this by introducing a set of control models $C=\{c_{1},\ldots,c_{m}\}$ chosen specifically for $f$. These control models are selected to match $f$’s domain, purpose or capabilities as closely as possible, while being clearly not ancestors of $g$. They are selected from publicly available models that share similar capabilities with $f$ but have been independently developed using different training datasets, architectures, or approaches. Such suitable control models are readily available in existing model repositories with sufficient variability in their development approaches. For instance, if $f$ is a French-language model for generating Python code, appropriate control models would include other French language models, Python code generation models, or ideally other French Python code generation models. On the other hand, if $f$ is a more general LLM such as Llama-3.2-1B, then the control models would include other general LLMs, such as Llama-3.2-3B, GPT2-large and Qwen2-1.5B. We do not have knowledge and influence beyond the domain or capabilities (e.g., “general” vs. “’French”). The number $m$ of control models depends on availability; generally, more control models from $f$’s domain lead to better baseline estimation. We observe empirically that there is enough diversity in existing pre-trained models such that constructing the control set does not require significant computational effort (e.g., additional training) or specialized machine learning expertise. Moreover, if the domain of the parent model $f$ is unknown, one can deduce its domain by manually analyzing the outputs of $f$ on some sample prompts. If such manual analysis is not feasible, we can just consider a larger and more diverse set of LLMs as the control set. For each control model $c_{i}$, we compute its similarity ratio $\mu_{i}$ with $g$ using the same process as before. These $\mu_{i}$ values form our baseline - they tell us how much $g$ typically agrees with models that operate in the same space as $f$ but are definitely not its ancestors. If the similarity $\mu$ between $g$ and $f$ significantly exceeds these baseline similarities $\mu_{i}$, this suggests a provenance relationship.
@@ 3.2 Model Provenance Tester
where $H_{0}^{i}$ is the null hypothesis that the similarity between $f$ and $g$ is less than or equal to the similarity between $c_{i}$ and $g$, and $H_{1}^{i}$ is the alternative hypothesis that the similarity between $f$ and $g$ is greater. To test each of the hypothesis $H^{i}$, we employ a z-test, which is a standard statistical test well-suited for comparing proportions like our similarity ratios when working with large samples. The z-test helps us determine if the observed difference between the two proportions ($\mu$ and $\mu_{i}$) is large enough to be statistically significant, or if it could have occurred by chance. The z-test produces a $p$-value, which represents the probability of observing such a difference in proportions if there were truly no difference between the models (i.e., if the null hypothesis were true). A small $p$-value (typically less than a predetermined significance level $\alpha$) suggests that the observed difference is unlikely to have occurred by chance, indicating that the similarity $\mu$ between $f$ and $g$ is indeed significantly larger than the similarity $\mu_{i}$ between $g$ and the control model $c_{i}$. We want all hypothesis tests to yield small $p$-values, indicating that the similarity $\mu$ is significantly higher than every baseline similarity $\mu_{i}$.
The significance level $\alpha$ represents our tolerance for incorrectly concluding that a difference is significant when it actually occurred by chance (a false positive). This threshold is set to a commonly used default value of $0.05$, meaning we accept a $5\%$ risk of claiming a significant difference where none truly exists. In our context, this would mean wrongly concluding that $\mu$ is significantly higher than some $\mu_{i}$ when the observed difference is merely due to chance.
@@
When conducting multiple hypothesis tests simultaneously, we want to maintain this same overall risk level of $\alpha$, regardless of how many tests we perform. However, running multiple tests increases our chances of obtaining at least one false positive across all tests (known as the family-wise error rate FWER). To control this cumulative risk, we employ the Holm-Bonferroni method [19], which adjusts the significance thresholds $\alpha_{k}$ for individual tests $H^{k}$ to ensure the overall false positive rate remains at or below our desired level of $\alpha$. The Holm-Bonferroni procedure works as follows. We first sort the individual $p$-values in ascending order: $p_{(1)}\leq p_{(2)}\leq\dots\leq p_{(m)}$. We then compare each p-value $p_{(k)}$ to its adjusted significance level $\alpha_{k}=\alpha/(m-k+1)$. Starting with the smallest $p$-value, we sequentially test each hypothesis. If a $p$-value $p_{(k)}$ is less than or equal to its corresponding $\alpha_{k}$, we reject the null hypothesis $H_{0}^{(k)}$ and proceed to the next one. This process continues until we encounter a $p$-value that exceeds its adjusted significance level, at which point we fail to reject the remaining null hypotheses. To conclude that $\mu$ is significantly greater than all $\mu_{i}$, we need to reject all null hypotheses $H_{0}^{i}$. By carefully controlling the FWER and ensuring that all null hypotheses are rejected through this method, we provide strong statistical evidence that the similarity between $f$ and $g$ is higher than that between $g$ and any control model $c_{i}$. This supports the assertion that $(f,g)$ has significantly higher similarity and thus constitutes a model provenance pair.
Algorithm 1 Model Provenance Tester for Pair $(f,g)$
$x_{1},\ldots,x_{T}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$triangleright Sample T prompts
$\mu\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(f(x_{j})=g(x_{j}))$triangleright Calc sim of f and g
for $i\leftarrow 1$ to $m$ do
     $\mu_{i}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(c_{i}(x_{j})=g(x_{j}))$triangleright Calc sim of $c_{i}$ and $g$
     $p_{i}\leftarrow\text{ZTest}(\mu,\mu_{i},T)$triangleright Obtain p-values
end for
$(p_{(1)},\ldots,p_{(m)})\leftarrow\text{Sort}(p_{1},\ldots,p_{m})$triangleright Sort p-values
for $k\leftarrow 1$ to $m$ do
     $\alpha_{k}\leftarrow\alpha/(m-k+1)$triangleright Holm-Bonferroni adjustment
     end if
end forreturn True $\triangleright$ Is a provenance pair
A pseudo-code of the tester is given in Algorithm 1. In summary, when the procedure returns True, it ensures that the total family-wise error rate (FWER) is controlled at the significance level of $\alpha=0.05$ or lower. This means we can confidently state that the similarity between $f$ and $g$ is significantly higher than between $g$ and any control model, supporting the existence of a provenance relationship. Conversely, when the procedure returns False, it indicates that we could not establish this higher similarity with the desired level of statistical significance. This may occur either because there is genuinely no significant similarity indicative of provenance or because the test lacked sufficient power under the given parameters (e.g., sample size or number of prompts) to detect it.
@@
$x_{1},\ldots,x_{T}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$triangleright Sample T prompts
@@
     $\mu_{i}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(f_{i}(x_{j})=g(x_{j}))$triangleright Calc sim of candidates
@@
     $\mu^{\prime}_{i}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(c_{i}(x_{j})=g(x_{j}))$triangleright Calc sim of controls

### v2

@@ Abstract
Large language models are increasingly customized through fine-tuning and other adaptations, creating challenges in enforcing licensing terms and managing downstream impacts such as protecting intellectual property or identifying vulnerabilities. We address this challenge by developing a framework for testing model provenance. Our approach is based on the key observation that real-world model derivations preserve significant similarities in model outputs that can be detected through statistical analysis. Using only black-box access to models, we employ multiple hypothesis testing to compare model similarities against a baseline established by unrelated models. On two comprehensive real-world benchmarks spanning models from 30M to 4B parameters and comprising over $600$ models, our tester achieves 909590-95% precision and 809080-90% recall in identifying derived models. These results demonstrate the viability of systematic provenance verification in production environments even when only API access is available.
@@ 1 Introduction
Platforms such as Amazon SageMaker and Hugging Face have enabled wide scale distribution of ML models, most notably large language models (LLMs) roziere2023code ; wu2024pmc-huggingface ; Intelligence2024 . The increase of publicly available foundation models and datasets, however, has also triggered concerns over unauthorized use of intellectual property (IP) and concerns of compromised datasets carlini2024poisoning or models hubinger2024sleeper . These issues are present not just in open-source ecosystems, but also for proprietary models that are hidden behind APIs anil2023palm ; chatgpt-api . For instance, concerns about model stealing attacks wherein one can extract the model parameters even for production-level models are on the rise carlini2024stealing ; tramer2016stealing ; orekondy2019knockoff . Similarly, there is growing concern that models may contain backdoors or vulnerabilities, making them susceptible to jailbreaking hubinger2024sleeper ; anil2024many ; zou2023universal . Despite best efforts to create a safe environment for the development of foundation models, there have already been instances of reported misuse china2024llm ; payload-malicious ; childabuse-diffusion .
This landscape highlights the growing need for model provenance testing. The problem of model provenance testing is as follows: Determine whether a target model has been derived from a foundational model by customizations such as fine-tuning. This problem has applications in tracking reuse of models not just in open marketplaces but also across product teams in large organizations. When a security or privacy audit finds a problem with a foundational model, it becomes important to identify which other models in use by the organization may be derived from the problematic one and take appropriate remedial actions (e.g. revoke, retrain, or fortify) to mitigate the risk of further non-compliant use. Model provenance tracking is often useful after the fine-tuned model has been deployed, and when authentic ground truth is unavailable or unreliable.
@@
Contributions..  In this paper, we design the first practical model provenance tester for LLMs that requires only query access. Our proposed techniques stem from a key empirical observation: The output distribution of fine-tuned models is often close to that of their parent model. This distance between a model and its true parent is typically smaller than that between the model and other unrelated models, making it possible to reliably trace back a derived model to the original parent. In order to keep assumptions to a minimum, we propose to employ the principled framework of statistical hypothesis testing. Specifically, we use black-box sampling and estimation to determine whether the distribution captured by the given model is close to that of the candidate parent. Such estimation can provide formal statistical significance measures, which can be used to check for the null hypothesis, i.e., the customized LLM is not close to the given parent model. Our approach is agnostic to the chosen metric for closeness and returns with the guarantee that the false positive rate is less than or equal to a user-specified threshold. We conduct an extensive empirical evaluation across two comprehensive benchmarks comprising over $600$ models from Hugging Face, ranging from 3030M to 44B model parameters and across diverse domains. Our tester achieves 9090-9595% precision and 8080-9090% recall in detecting model provenance, even with a limited number of queries.
2 Model Provenance Testing
Pretraining LLMs involves significant investment, requiring substantial computational resources costing millions of dollars in infrastructure and thousands of GPU hours. When Company A releases a pretrained LLM denoted as $f$, it employs specific licensing terms crucial for protecting this investment, maintaining competitive advantage, and controlling the model’s usage chatgpt-api ; anil2023palm . Startup B might download $f$, perform only fine-tuning or other light customization (mixture-of-experts, prompt engineering), but claim to have pretrained their model $g$ from scratch, thereby circumventing licensing requirements and misrepresenting company A’s work. In such cases, we want to be able to determine if $g$ is derived through fine-tuning of $f$ and resolve the model provenance problem.
@@ 2 Model Provenance Testing
We consider model provenance testing framework with minimal assumptions only, where the tester can query the models $f,g$ on arbitrary prompts (e.g., through an API) and get responses. The tester has no access to the training datasets, embeddings, hyperparameters used by either company, or any information about potential modifications performed by Startup B. This mirrors real-world conditions where companies do not always disclose their training procedures, data sources, or modification techniques, making the provenance testing problem both practical and challenging.
Our proposed methods are evaluated for a non-adaptive adversary, i.e., the Startup B is not aware of the strategy of the model provenance tester. On the other hand, it is not clear if more advanced techniques for evasion of provenance detection deployed by adaptive adversaries can provide real benefit due to the randomness used in our tester (sampling from infinite set of prompts) and the fact that adversarial fine-tuning may come at the cost of reduced model performance on intended task.
2.1 Related Work
The problem of testing provenance is similar to designing schemes to prove ownership of a model. We provide more details on related work in Appendix A and summarize main approaches here.
Fingerprinting.  Fingerprinting aims to create an identifier of the model for different downstream purposes such as IP protection. Most prior work require white-box access or additional knowledge such as intermediate LLM modules zeng2024huref , access to training and testing datasets and changes to the fine-tuning xu2024instructional . One type of approaches detect copies or tampering by hashing the model weights chen2023perceptual ; xiong2022neural . Other works aim to create fingerprints by estimating the decision boundary via adversarial examples lukas2019deep ; xu2024united ; peng2022fingerprinting . However, adversarial-based approaches can still have a high false positive rate, i.e., two unrelated models have the same fingerprinting samples cao2021ipguard . Follow-up work generates universal perturbations that can be added to samples to match the target model lukas2019deep . This can be prone to adversarial evasion huang2023can . Thus, recent work constructs a set of adversarial examples that are connected with each other and that characterize the decision boundary more robustly xu2024united . In addition to model fingerprinting via adversarial perturbations, there are works that focus on measuring the model similarity huang2023can ; guan2022you ; chen2022copy ; jia2021zest ; dong2023rai2 . For example, one work selects samples with different prediction results across two groups of reference models guan2022you . This approach utilizes their pairwise correlation to identify target stolen image classification models. Other approaches also rely on the internal representations to compare models, without focusing on provenance. For instance, one work locally approximates the models under test with linear models and then computes the cosine distance between the weights of the linear model jia2021zest . One recent work in black-box access requires specialized prompting and an additional model’s embeddings to identify models pasquini2024llmmap . We highlight that these approaches do not focus on the question of model provenance as fingerprints may not be robust to fine-tuning and other customizations. Our tester also has minimal assumptions of black-box access to only next-token predictions. While our work considers this measure of similarity, our approach returns with a bounded false positive rate, which is metric-agnostic.
Watermarking.  Other approaches prove ownership via embedding watermarks in the model weights chen2019deepmarks ; wang2021riga ; uchida2017embedding , hidden-layer activations darvish2019deepsigns or other model features li2022defending ; chen2021you . Another type of watermarks are achieved by learning triggers or backdoors that produce predefined outputs jia2021entangled ; shao2024explanation ; yan2023rethinking ; peng2023you ; cao2021ipguard ; zhang2018protecting . These rely on implementing changes to the white-box model, fine-tuning or training the model on a specialized dataset and can affect the model’s performance. Specialized model ownership schemes have also been proposed for other models such as graph neural networks zhou2024revisiting ; waheed2024grove . Most existing approaches do not have provable guarantees that model ownership can be verified with a given confidence, so the verification is often empirically determined. It is particularly challenging to preserve the watermark detection after applying model customizations krauss2024clearstamp .
The focus of prior work has not been on designing tests for determining provenance under model customizations and do not work under the same minimal assumptions. For instance, recent work proposed a detection framework for fine-tuning and pruning that still requires white-box access to the victim model to generate test cases and it does not consider LLMs chen2022copy . Moreover, our benchmarks are much more extensive with hundreds of models under diverse customizations.
3 Approach
Our approach to testing model provenance is based on a key observation: fine-tuning and other model derivation techniques typically result in only limited changes to the original model, as they primarily adapt the model for new tasks. After fine-tuning, the derived model $g$ may remain similar to its parent model $f$, as the process focuses on refining specific capabilities rather than creating fundamental changes to the model distribution.
3.1 Model Provenance Tester
Since the tester has only query access to the models, the only data it can collect is from providing inputs (called prompts) and analyzing the corresponding output tokens. Furthermore, due to our minimal assumptions and lack of information about the training datasets, the tester queries the models on randomly chosen short sequences, further denoted as a prompt space $\Omega$.
The tester independently samples from $\Omega$ a set of $T$ prompts $x_{1},\ldots,x_{T}$, queries each model on the same set of prompts, and then compares their output tokens pairwise. For each prompt, we compare only the first output token generated by each model; however, $n$-grams could also be considered. The similarity between two models $f$ and $g$ is then calculated as the proportion $\mu$ of prompts on which the models produce the same output token: $\mu=\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(f(x_{j})=g(x_{j}))$.
@@ 3.1 Model Provenance Tester
However, the ratio $\mu$ by itself may not be sufficient to determine whether $f$ and $g$ are similar, since even two unrelated models might agree on some proportion of outputs by chance. To assess the significance of $\mu$, we introduce a set of control models $C=\{c_{1},\ldots,c_{m}\}$ for the model $f$ and compare their similarities to $g$. These control models are sampled from the set of available LLMs with distinct architectures or training data, to ensure diversity. Our aim is to compare $\mu$ to the similarity ratios $\mu_{i}$ between each control model $c_{i}$ and $g$. The number $m$ of control models and their sampling strategy depend on availability; generally, the larger the set of control models, the better the estimation of similarity among unrelated models.
@@
where $H_{0}^{i}$ is the null hypothesis that the similarity between $f$ and $g$ is less than or equal to the similarity between $c_{i}$ and $g$, and $H_{1}^{i}$ is the alternative hypothesis that the similarity between $f$ and $g$ is greater. To test each of the hypothesis $H^{i}$, we employ a z-test, which is a standard statistical test well-suited for comparing proportions like our similarity ratios when working with large samples. The z-test produces a $p$-value, which represents the probability of observing such a difference in proportions if there were truly no difference between the models (i.e., if the null hypothesis were true). A small $p$-value (typically less than a predetermined significance level $\alpha$) indicates that the similarity $\mu$ between $f$ and $g$ is indeed significantly larger than the similarity $\mu_{i}$ between $g$ and $c_{i}$.
@@
When conducting multiple hypothesis tests simultaneously, we want to maintain this same overall risk level of $\alpha$, regardless of how many tests we perform. However, running multiple tests increases our chances of obtaining at least one false positive across all tests, known as the family-wise error rate (FWER). To control this cumulative risk, we employ the Holm-Bonferroni method holm1979simple , which adjusts the significance thresholds $\alpha_{k}$ for individual tests $H^{k}$ to ensure the overall false positive rate remains at or below our desired level of $\alpha$. We illustrate this basic tester idea in Figure 1.
@@
$x_{1},\ldots,x_{T}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$ $\triangleright$ Sample T prompts
@@
  $\mu_{i}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(f_{i}(x_{j})=g(x_{j}))$ $\triangleright$ Calc sim of candidates
@@
  $\mu^{\prime}_{i}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(c_{i}(x_{j})=g(x_{j}))$ $\triangleright$ Calc sim of controls